OpenStudy (anonymous):

solve 2/3x + 5 = 1 ....is the answer x = -8/3?

OpenStudy (anonymous):

is the first term (2/3) times x or 2 divided by 3x?

OpenStudy (anonymous):

2 divided by 3x

OpenStudy (anonymous):

sorry so long - i was called away

OpenStudy (anonymous):

if you substitte -8/3 for x in the equation: 2 / (3* -8/3) + 1 = 2 / -8 + 1 = (-1/4) + 1 = 3/4 so no thats is not the answer

OpenStudy (anonymous):

would u like me to work thru it ?

OpenStudy (anonymous):

2/3x + 5 = 1 x thru by 3x 2 + 15x = 3x 2 = -12x x = -1/6

OpenStudy (anonymous):

ohh ok thanks for the help

OpenStudy (anonymous):

no probs - my grandkids kept interrupting me!!!